OpenStudy (anonymous):

kindly help me to factor this: 3x^2y^4-27

OpenStudy (lalaly):

3(xy^2-3)(xy^2 +3)

OpenStudy (anonymous):

=3((xy^2)^2-3^2) =3(xy^2+3)(xy^2-3)

OpenStudy (anonymous):

what factoring technique did you used?

OpenStudy (anonymous):

Difference of squares

OpenStudy (anonymous):

took 3 out as a factor then the given equation becomes similar to the identity \[(a^2-b^2)=(a+b)(a-b)\]